### 24 Jan 2024 v2

@@ 4 Discussion
Table 8: Causal Links for the IRA case. We show the value of the maximal causal effect, $I_{i\to j}^{{\mathrm{CE,max}}}={\mathrm{max}}_{0<\tau\leq\tau_{{\mathrm{max}}}}\left|{I_{i\to j}^{{\mathrm{CE}}}(\tau)}\right|$ between each pair (i, j) of activity time series, where $\tau_{max}=18\times 15$min=4.5h is the maximal time lag considered, with standard errors. The arrows indicate the the direction of the causal effect. For each activity time series, we indicate in bold the most important drivers of activity (excluding themselves). In blue, we highlight the auto-correlation of each node.
@@
Table 9: Causal Links for the IRA+S case. We show the value of the maximal causal effect, $I_{i\to j}^{{\mathrm{CE,max}}}={\mathrm{max}}_{0<\tau\leq\tau_{{\mathrm{max}}}}\left|{I_{i\to j}^{{\mathrm{CE}}}(\tau)}\right|$ between each pair (i, j) of activity time series, where $\tau_{max}=18\times 15$min=4.5h is the maximal time lag considered, with standard errors. The arrows indicate the the direction of the causal effect. For each activity time series, we indicate in bold the three most important drivers of activity (excluding themselves). In blue, we highlight the auto-correlation of each node.

### 13 Mar 2024 v3

@@ Acknowledgements
Table 6: Causal Links: IRA. We show the value of the maximal causal effect, $I_{i\to j}^{{\mathrm{CE,max}}}={\mathrm{max}}_{0<\tau\leq\tau_{{\mathrm{max}}}}\left|{I_{i\to j}^{{\mathrm{CE}}}(\tau)}\right|$ between each pair (i, j) of activity time series, where $\tau_{max}=18\times 15$min=4.5h is the maximal time lag considered, with standard errors. The arrows indicate the the direction of the causal effect. For each activity time series, we indicate in bold the most important drivers of activity (excluding themselves). In blue, we highlight the auto-correlation of each node.
@@
Table 7: Causal Links: Suspended. We show the value of the maximal causal effect, $I_{i\to j}^{{\mathrm{CE,max}}}={\mathrm{max}}_{0<\tau\leq\tau_{{\mathrm{max}}}}\left|{I_{i\to j}^{{\mathrm{CE}}}(\tau)}\right|$ between each pair (i, j) of activity time series, where $\tau_{max}=18\times 15$min=4.5h is the maximal time lag considered, with standard errors. The arrows indicate the the direction of the causal effect. For each activity time series, we indicate in bold the three most important drivers of activity (excluding themselves). In blue, we highlight the auto-correlation of each node.